OpenStudy (lovelyanna):

Write an equation in slope-intercept form for the line passing through (–1, 10) and (5, 8). A. y = 1/3x + 29/3 B. y = -1/3x - 29/3 C. y = -1/3x + 29/3 D. y = 1/3x - 29/3 I know the answer is C, but can someone please explain? Thanks

OpenStudy (lovelyanna):

@math&ing001

OpenStudy (anonymous):

slope= 8-10=-2 over 5--1 which is plus so equals 6. slope is -2/6 or -1/3

OpenStudy (lovelyanna):

I know the slope, can you explain the rest?

OpenStudy (anonymous):

y-1=-1/3(x-10) a negative times a negative is positive so thats where the +29 would come from

OpenStudy (lovelyanna):

and the x by -1/3. And the 3 in 29/3. How about those?

OpenStudy (anonymous):

wait thats not righ lol

OpenStudy (lovelyanna):

What isn't right?

OpenStudy (anonymous):

what i said wasnt right, im trying to figure out how we got to 29/3

OpenStudy (lovelyanna):

Ok, but what is x?

OpenStudy (anonymous):

the x is part of point slope form

OpenStudy (anonymous):

y - y1 = m(x - x1)

OpenStudy (lovelyanna):

So... still confused :/

OpenStudy (anonymous):

If we filled it in it would be y-10=-1/3(x+1)

OpenStudy (lovelyanna):

Oh, ok. How about 29/3?

OpenStudy (anonymous):

ok 29/3 =9.66 so when we distribute the -1/3 over to +1 we get -1/3 again. now we add 10 to the other side to get in numerical terms 9.66

OpenStudy (lovelyanna):

I still don't get how you got 9.66.. I'm sorry, can you explain more, this is very confusing.. :(

OpenStudy (anonymous):

y-10=-1/3(x+1) --> distribute -1/3 to x and 1 to get y-10=-1/3x - 1/3. Now add 10 to the other side. 10+-1/3 essentially equals 9.66

OpenStudy (anonymous):

9.66 is 29 over 3

OpenStudy (lovelyanna):

Oh, that makes much more sense! thanks

OpenStudy (anonymous):

do yu get it? I dont know how to get the 29/3 sorry

OpenStudy (lovelyanna):

But you just got it.. 9.66, right?

OpenStudy (lovelyanna):

@math&ing001 can you help?

OpenStudy (anonymous):

yes 9.66 equals 29/3

OpenStudy (anonymous):

@campbell_st can you help aswell?

OpenStudy (math&ing001):

Gemme a sec :)

OpenStudy (math&ing001):

Ok so you know how to find the slope, I'll just explain the rest. Formula for the slope gives: \[m=-\frac{ 1 }{ 3 }=\frac{ y-8 }{ x-5 }\] So: \[y-8=-\frac{ 1 }{ 3 }x+\frac{ 5 }{ 3 }\] And finally: \[y=-\frac{ 1 }{ 3 }x+\frac{ 29 }{ 3 }\]

OpenStudy (anonymous):

Thank you :)

OpenStudy (math&ing001):

Welcome !

OpenStudy (anonymous):

one question how did you go from 8 to 29

OpenStudy (lovelyanna):

Ok, thank you both :)

OpenStudy (math&ing001):

Add 8 on both sides, then 8=24/3 so (24+5)/3=29/3. yw @LovelyAnna

OpenStudy (anonymous):

ohhh ok I see thanks again

OpenStudy (lovelyanna):

@jigglypuff314

OpenStudy (lovelyanna):

Hi Emily! Can you explain? Maybe I can understand from you a bit better :)

jigglypuff314 (jigglypuff314):

slope-intercept form: y = mx + b when m = slope and b = y-intercept so first find the slope between (–1, 10) and (5, 8) \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]when given points (x1, y1) and (x2, y2)

OpenStudy (lovelyanna):

I know the slope is -1/3, how about x and 29/3?

jigglypuff314 (jigglypuff314):

so you plug in m = -1/3 to get y = -1/3 x + b then you would temporarily plug in a point to get the value of b

OpenStudy (lovelyanna):

Mhm.. I understand it till there, how would you get b after that? I don't get that part..

jigglypuff314 (jigglypuff314):

so you plug in m = -1/3 to get y = -1/3 x + b then you would temporarily plug in a point to get the value of b so if you plug in point (-1, 10) then you'll get 10 = (-1/3)(-1) + b solve for b

OpenStudy (lovelyanna):

Hmm.. confusing :/

jigglypuff314 (jigglypuff314):

what part do you not get? :)

OpenStudy (lovelyanna):

How to solve for b

jigglypuff314 (jigglypuff314):

so if you plug in point (-1, 10) then you'll get 10 = (-1/3)(-1) + b negative times negative equals positive so 10 = 1/3 + b then subtract 1/3 from both sides b = 10 - 1/3 10 = 30/3 \[so~~~b =\frac{ 30 }{ 3 } - \frac{ 1 }{ 3 }\]therefore b = 29/3

OpenStudy (lovelyanna):

Oh, I see! thanks :)

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (lovelyanna):

There is another question I got wrong.. can you check it?

OpenStudy (lovelyanna):

Select the pair of equations whose graphs are perpendicular. A. 2x – 8y = 9 12x – 3y = 7 B. 2y = –3x + 5 2x + 3y = 4 C. 3x + 6y = 8 y = 2x – 8 D. y = 2x – 7 x + y = 3 I choose B, but that isn't correct.. maybe C?

jigglypuff314 (jigglypuff314):

yeah C :) because perpendicular slope is the flipped opposite

OpenStudy (lovelyanna):

Thought so :) can you show me its equation?

jigglypuff314 (jigglypuff314):

what equation? 3x + 6y = 8 -> y = -1/2 x + 4/3 y = 2x – 8 -1/2 is the flipped opposite of 2

OpenStudy (lovelyanna):

I mean how you could solve the question

jigglypuff314 (jigglypuff314):

solve the system of equations of ? 3x + 6y = 8 -> y = -1/2 x + 4/3 y = 2x – 8 y = y so -1/2 x + 4/3 = 2x - 8 4/3 + 8 = 2x + 1/2 x 28/3 = 5/2 x x = 56/15

OpenStudy (lovelyanna):

Oooh, I get it! :D Thanks!

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (lovelyanna):

There is 1 last question I want to make sure of, you think you can help?

jigglypuff314 (jigglypuff314):

sure :)

jigglypuff314 (jigglypuff314):

2t + 18 ≥ 14 2t + 18 < 20 2t ≥ -4 2t < 2 t ≥ -2 t < 1 so t is between -2 and 1 with closed hole at -2 and open hole at 1

OpenStudy (lovelyanna):

I don't get that.. but this is how I solved it: 14 ≤ 2t + 18 < 20 subtract 18 from both sides: -4 ≤ 2t < 2 divide 2 by both sides -2 ≤ t < 1 Is this correct?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (lovelyanna):

Alright, thanks Emily! =) I really appreciate your help!

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (lovelyanna):

Wish I could help you.. but your in a higher grade than me :(

jigglypuff314 (jigglypuff314):

lol, I just got a whole lot of stuff that I need help on xD It's fine though :P I like helping people out with simple math <3

OpenStudy (lovelyanna):

Oh.. sorry :( If you say so.. that's what I like about you.. your soo helpful, and kind!

jigglypuff314 (jigglypuff314):

I try to be :P

OpenStudy (lovelyanna):

Btw.. which equation is better for the last question, mine or yours?

OpenStudy (lovelyanna):

You already are! :)

jigglypuff314 (jigglypuff314):

yours :P and there are a few people on here that I really don't like helping >,<

OpenStudy (lovelyanna):

You think so? I wonder why.. O.o Really, like who? oh.. you mean people who use you to just answer there questions? :P

jigglypuff314 (jigglypuff314):

yeah... like they figured out the perfect way to just get answers out of me >,< like posting a whole lot of questions at once, or guilt tripping me cuz they need to pass with a certain score within a certain amount of time... I'm not gonna name names tho :P

OpenStudy (lovelyanna):

Well, I post a lot of questions in one place.. you think I'm one of them? :O eeek!!

jigglypuff314 (jigglypuff314):

nope, cuz most of the time you're checking your answers :) gimme a sec, I'll find an example...

OpenStudy (lovelyanna):

Oh, ok.. if you say so.. phew! :) Ok, I'll wait

jigglypuff314 (jigglypuff314):

http://openstudy.com/study#/updates/52b348bde4b03e81ff954681

OpenStudy (lovelyanna):

OMG.. :O That is NOT cool!

jigglypuff314 (jigglypuff314):

indeed -.-

OpenStudy (lovelyanna):

Hmm.. your too nice to have answered them >,< cheaters..

jigglypuff314 (jigglypuff314):

there's a lot of cheaters on this site -.-" I'm thinking about signing up to try to become an ambassador on this site...

OpenStudy (lovelyanna):

What is an ambassador? And can you tell me the difference between a solid circle and an empty one? Does ≤ and < have anything to do with it? http://static.k12.com/calms_media/media/212500_213000/212713/1/556fb09d1057b44d43dd6c45f65425352134d2b7/ALGTests_L2_S1_U9_Q9_B.gif

jigglypuff314 (jigglypuff314):

an ambassador are the people on this site that have the purple A next to their name, they get to welcome newcomers and have a chance to become a moderator :P solid hole means would refer to the "or equal to" part if it doesn't have that part then it would be an empty one :)

OpenStudy (lovelyanna):

Oh, thats cool! How will you become one? I get is, so empty = < or > And solid = \[\le or \ge \]

jigglypuff314 (jigglypuff314):

http://blog.openstudy.com/2013/09/09/the-openstudy-ambassador-program/ and yep :)

jigglypuff314 (jigglypuff314):

bye and good luck :)

OpenStudy (lovelyanna):

@jigglypuff314

jigglypuff314 (jigglypuff314):

hi :)

OpenStudy (lovelyanna):

Hi, I'm back :) I was wondering if you can help me with drawing graphs?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (lovelyanna):

Thanks :) These: 2y = –3x + 5 2x + 3y = 4 3x + 6y = 8 y = 2x – 8 I'm not soo good with graphs.. :(

jigglypuff314 (jigglypuff314):

well you could convert it into slope intercept form then you know how to graph stuff when they're in slope intercept form right? :)

OpenStudy (lovelyanna):

Hmm.. give me a min

OpenStudy (lovelyanna):

I don't know.. :'(

jigglypuff314 (jigglypuff314):

2y = –3x + 5 -> y = -3/2 x + 5/2 2x + 3y = 4 -> y = -2/3 x + 4/3

OpenStudy (lovelyanna):

Woah.. how did you get all that? O.o

jigglypuff314 (jigglypuff314):

2y = –3x + 5 divide both sides by 2 y = -3/2 x + 5/2 2x + 3y = 4 3y = -2x + 4 divide both sides by 3 y = -2/3 x + 4/3

OpenStudy (lovelyanna):

Oh, I see. Now we go to the graphs?

jigglypuff314 (jigglypuff314):

yep :)